OpenStudy (18jonea):

Solve the equation a2 + 4a = -3 by graphing or by factoring. a = 1, a = 3 a = -1, a = 3 a = -3, a = 1 a = -3, a = -1 @robtobey

OpenStudy (18jonea):

@Michele_Laino

Directrix (directrix):

Try factoring: a2 + 4a = -3 a^2 + 4a + 3 = 0 @18jonea Factor the expression on the left side of the equation.

OpenStudy (18jonea):

isthe answer d

Directrix (directrix):

a^2 + 4a + 3 = 0 (a + 3) * (a + 1) = 0 a + 3 = 0 or a + 1 = 0 a + ? or a = ?

Directrix (directrix):

>>isthe answer d Correct.